OpenStudy (anonymous):

I need some assistance with this word problem please. The length of a rectangle is 3yd more than twice its width, and the area of the rectangle is 35^2. Find the dimensions of the rectangle.

OpenStudy (rsadhvika):

let length = l width = w

OpenStudy (rsadhvika):

The length of a rectangle is 3yd more than twice its width => l = 2w + 3

OpenStudy (rsadhvika):

area of the rectangle is 35 => area = l*w => 35 = l * w 35 = (2w+3)w 35 = 2w^2 + 3w 2w^2+3w-35 = 0

OpenStudy (rsadhvika):

you can solve the quadratic equation for w ?

OpenStudy (anonymous):

i think so ..I am so not sure of myself anymore :(

OpenStudy (rsadhvika):

hmm

OpenStudy (anonymous):

-35+3w+2w^2 +0 (-5-1)(7-2)=0 Is this right so far?

OpenStudy (anonymous):

(-5-1w)(7-2w) i mean

OpenStudy (rsadhvika):

yes !!

OpenStudy (rsadhvika):

so w = 7/2

OpenStudy (rsadhvika):

find what l is

OpenStudy (anonymous):

3.5

OpenStudy (anonymous):

would be the decimal right?

OpenStudy (rsadhvika):

yes 3.5 is good :P

OpenStudy (anonymous):

L would be 10?

OpenStudy (anonymous):

cause you have the 2(3.5) +3 +10

OpenStudy (rsadhvika):

correct ! very good work :)

OpenStudy (anonymous):

thank you! you are an awesome teacher

OpenStudy (rsadhvika):

np... yw:))